# Rubber Friction: Theory, Mechanisms, and Challenges

### v2

@@ 9 Rubber friction on ice and snow
Friction on ice is a fascinating and ancient subject [87, 88, 89, 90, 91], with pioneering work by Faraday [92] over 150 years ago, in which he proposed that the surface of ice is covered by a liquid-like film of water.
This phenomenon is now known as premelting. The presence of a premelted layer depends on both temperature and the hydrophobic or hydrophilic nature of the counter surface [93, 94]. For inert hydrophobic materials, such as polytetrafluoroethylene (Teflon), such a layer may appear well below the ice melting point. In contrast, for hydrophilic surfaces such as silica, a pre-melted water film may not occur at all or only very close to the ice melting point. Here we also note a recent study by Atila, Sukhomlinov and Müser who used molecular dynamics and showed that ice surfaces can “liquefy” without melting thermodynamically by cold displacement-driven amorphization. However, this study also showed that for the friction to be low the countersurface must be hydrophobic, or else the sliding speed so high that water is produced by frictional heating[95].
Another explanation for the formation of a water film on ice was proposed by Thomson [96], who suggested that it results from pressure-induced melting. However, this mechanism is likely invalid, as plastic deformation of the ice usually occurs before the pressure is sufficient to induce melting.
The importance of frictional heating was proposed by Bowden and Hughes [97], who argued that frictional heating causes surface melting of the ice if the sliding speed is large enough. This mechanism has been confirmed by experimental studies [98]. The low friction observed on ice surfaces can therefore be attributed to a combination of premelting, cold displacement-driven amorphization and frictional heating effects [93]. One very important observation for friction on ice is that when the sliding speed increases the friction drops, which imply smaller frictional heating. This will shift to higher sliding speeds the onset of ice melting, and can explain the velocity dependency of the friction on ice[99].
@@
  • At temperature very close to the ice bulk melting temperatures, or at high sliding speeds, a water (or water-like) film is present at the interface, and the friction is dominated by viscoelastic deformation of the rubber by the ice roughness[84, 93, 100].
@@
  • At lower temperatures and lower sliding speeds, the water layer is absent. In this case, the friction arises from both the viscoelastic response of the rubber and the shear in the rubber-ice contact regions, or between ice fragments attached to the rubber surface and the underlying ice, which constitutes an ice–ice interaction [93].
@@
  • At even lower temperatures (below $T_{\rm g}$), the viscoelastic contribution becomes negligible. As the rubber becomes stiff, it can cause wear of the ice asperities upon contact, leading to increased friction due to a plowing-type process [100, 101].
@@
In Ref. [84], the Persson rubber friction theory was applied to rubber sliding on ice. The study incorporating both the adhesive contribution from the real area of contact and the viscoelastic contribution from the rubber. A temperature-dependent large wavenumber cutoff, $q_{1}(T)$, was introduced to account for the smoothing of short-wavelength roughness by plastic deformation of the ice. Since the ice penetration hardness depends on the temperature and indentation speed dependent, the cut-off depends on the ice surface temperature and on the rubber sliding speeds. The model also includes the effect of frictional heating in the rubber, which leads to a shift in the viscoelastic modulus to higher frequencies. The model offers a consistent explanation of the measured friction behavior for three very different tread compounds.
@@
Figure 52: (a) The measured rubber-ice friction coefficient as a function of the logarithm of the sliding speed for several temperatures indicated. For low temperatures the friction is very similar to ice sliding on ice suggesting that the friction is due to ice wear particles on the rubber surface sliding on the ice surface. (b) Comparison between experimentally measured friction coefficients (symbols) and theoretical predictions of the viscoelastic contribution (solid lines), based on the theory developed in [84]. At the higher temperature (black), the measured friction closely matches the viscoelastic prediction, indicating that a thin premelted water film forms at the rubber–ice interface, and that friction is dominated by viscoelastic deformation. At the lower temperature (green), the measured friction is significantly higher than the theoretical prediction, suggesting additional contributions from interfacial shear. Adapted from Ref. [93].
@@
Figure 53: The friction coefficient as a function of the logarithm of the sliding speed for ice on ice at different temperatures. Adapted from Ref. [102].
@@
Fig. 52(a) shows the rubber friction on ice as a function of the sliding speed for several different temperatures. In Fig. 52(b) the experimental measurements are compared with theoretical predictions based on the viscoelastic contact mechanics theory developed in [84]. At the higher temperature ($T=-5.4^{\circ}{\rm C}$), the experimental data nearly coincide with the predicted viscoelastic contribution, indicating that a thin premelted water film likely forms at the rubber–ice interface, and that friction is dominated by viscoelastic deformation. In contrast, at the lower temperature ($T=-30^{\circ}{\rm C}$), the measured friction coefficient significantly exceeds the viscoelastic prediction. For $T<-20^{\circ}{\rm C}$ the friction of rubber on ice is similar to the friction of ice on ice (see Fig. 53), suggesting a substantial contribution from interfacial shear between ice fragments and the ice surface.
@@
Figure 54: (a) Rubber block squeezed in contact with an ice surface covered by stone particles. (b) During sliding ice plowing occurs which result in a friction force which is maximal when the projected area of the particles occupy the fraction $\psi\approx 2\sigma_{0}/\sigma_{\rm P}$ of the nominal surface area, where $\sigma_{0}$ is the nominal contact pressure squeezing the rubber block against the substrate and $\sigma_{\rm P}$ the ice (temperature-dependent) penetration hardness. The maximum plowing friction coefficient is $\mu\approx 1$. (c) The ice roughness with wavelength larger than the sand particles will give a viscoelastic contribution to the friction force.
@@
Suppose there are $N$ sand particles in the tire-road footprint. If the concentration of sand particles is high enough, the full normal load of the tire will act on the sand particles, giving the force $f_{\rm N}=F_{\rm N}/N$ that squeezes each sand particle against the ice surface. If the sand particles are spheres with radius $R$, and if $\sigma_{\rm P}$ is the ice penetration hardness, then the contact area projected onto the horizontal $xy$-plane is $A=f_{\rm N}/\sigma_{\rm P}$.
The maximum friction occurs when the force $f_{\rm N}$ is large enough to squeeze half of the particle into the ice. If $f_{\rm N}$ exceeds the value required to embed half of the sand particle, the entire particle may be pressed into the ice, and the contact would become rubber–ice contact. For the stationary case, the contact area projected onto the horizontal $xy$-plane would be $\pi R^{2}$, and the maximum penetration resistance is $\pi R^{2}\sigma_{\rm P}$. However, for the sliding (plowing) case, we assume the contact area is half of this, $A=\pi R^{2}/2$, since the sand particle only contacts the ice on one side. In this case, the ice wear track will have a width of $2R$, and the vertical cross-sectional area of the plowing track will be $A_{\rm plow}=\pi R^{2}/2$.
@@
If this distance is larger than $R$, the rubber may contact the ice, and part of the normal force (the load) would then be carried by direct rubber–ice contact. As a result, the force $f_{\rm N}$ acting on the sand particles would be reduced. Hence, for (38) to be valid, we must have
@@
Figure 55: Measured rubber friction coefficient as a function of temperature for rubber sliding on clean ice, on ice with a thin snow film, and on ice covered by a thin snow film and sand. The friction was measured using the British Pendulum Tester (sliding speed on the order of $\sim 1\ {\rm m/s}$). Adapted from [106]
@@
Sanding with a high concentration of sand particles $\psi>0.2$ may result in friction smaller than that on a clean surface, unless the temperature is close to the ice melting temperature (see Fig. 55). This is consistent with some experimental observations, where the friction level of rubber on sanded ice was found to be lower than that on clean ice [107, 108]. Norwegian airport maintenance personnel have also pointed out that in some cases sanding has a negative effect on friction level, and they commonly refer to it as a “ball-bearing” effect [106, 109]. (However, there is no general reason for the sand particles to roll on the ice surface.)
This behavior is expected when the sand particle concentration is high because in this case (44) predicts $\mu<0.2$ if $\psi>9(\sigma_{0}/\sigma_{\rm P})$, and using $\sigma_{0}=0.2\ {\rm MPa}$ and $\sigma_{\rm P}=10\ {\rm MPa}$ gives $\psi>0.18$. Since sanding removes the adhesive contribution to friction and the viscoelastic contribution from roughness with wavelengths shorter than the sand particles, this may result in a friction coefficient smaller than that observed on clean ice for $T<-5^{\circ}{\rm C}$.
Finally, we note that wear tracks have been observed even when rubber slides on clean ice surfaces [110, 111]. In Ref. [110], it was suggested that the narrow wear tracks observed may have resulted from contamination particles (dust), which cannot be avoided when performing experiments in the normal atmosphere. However, in Ref. [111], it was suggested that hard filler particles on the rubber surface scratch the ice surface and result in the strongly modified (smoothed) ice surface observed after repeated sliding on the same area. In addition, the stress in rubber asperities that are strongly deformed (strain of order unity) is on the order of the modulus $E_{*}$, which can be similar to the yield stress of ice. Hence, wear tracks may arise if large enough rubber asperities occur, even without considering the fillers.
@@
The friction of rubber on snow shares many similarities with that on ice, as snow can be regarded as loose packed ice particles. However, snow flakes can have a complex dendrite structure or consist of small nearly spherical ice grains and these two types of snow could have different frictional properties against rubber or other solids. Although dedicated studies on the frictional behavior of rubber on snow surfaces remain limited (mostly for tires) [112, 113, 114, 115], there are many studies of friction for polymer sliders (skis) on snow [116, 117, 118]. This field was pioneered by Bowden and Hughes [119], who used skis to investigate friction on snow and found similar characteristics to friction on ice, although they did not propose a specific model for snow friction. For stiff solid sliders like skis Colbeck [120, 121, 122, 123, 124] identified the key components of snow friction as plowing, direct solid-solid contact (i.e., adhesive friction), meltwater lubrication due to frictional heating, and capillary forces from liquid bridges in partially melted snow.
Some of these mechanisms for friction between snow and hard polymers may also apply to rubber materials, but additional considerations are needed due to their viscoelastic nature (see Fig. 56). Any fluctuation in the (compacted) snow layer height which is not flattened in the tire footprint could generate a viscoelastic contribution to the friction during slip. However, if the frictional shear stress is larger than the yield properties of the ice in shear, the the compacted snow may instead undergoes (internal) shear deformations, but this too would contribute to the friction force.
The shear strength of snow compressed by a normal stress $\sigma$ has been studied experimentally. Fig. 57 show results for the shear stress immediately after compression (no aging) as relevant for tire applications. The shear stress increases nearly linear with the applied pressure and is of similar magnitude as the normal stress. This imply that if the area sheared would be similar size as the nominal tire-road contact area a friction coefficient of order unity could be expected. In reality, the area where the shear plane is inside the compacted snow rather than at the rubber-snow interface may be much smaller than unity. However, if the tire would have a large concentration of cavities as in Fig. 58, and if the pressure between the snow and the rubber in the cavities would be similar as the nominal tire-road footprint elsewhere, then a big friction coefficient may be expected.
@@
Figure 58: If the normal pressure acting on the snow in the rectangular cavities on the rubber surface is similar to the nominal pressure acting in the tire-road footprint then shearing the snow-snow interface as indicated by the dashed line could give a contribution to the friction coefficient which, depending on the slip velocity, could be of order the fraction of the the nominal contact area covered by the cavities, e.g. 0.5 if the cavities cover $50\%$ of the total area.
Overall, the friction of rubber on snow depends not only on the material properties but also on the snow morphology and water content (which is temperature-dependent):
  • Wet snow or slush (snow mixed with water, salt, mud ..) can act as a high viscosity fluid which cannot be squeezed out from the tire-road footprint even at relative low speed. Hence it may act as a high viscosity lubricant which could result in (viscous) hydroplaning and low friction.
  • On dry snow friction arises in part by shearing snow contacts and in part by viscoelastic deformation of the rubber by inhomogenities in the height of the compressed snow film. There will also be a contribution from shearing the rubber-snow interface if slip occur at this interface.
For rubber tires, tread design and sipes play a critical role. Compared to smooth surfaces, treads with sipes generate additional plowing at the edges of the bent rubber segments. In addition compacted snow can enter the sipes which result in a snow-snow interface (of compacted snow) which is sheared during slip. These effects significantly enhance snow grip [112, 114, 115]. Tread depth and void ratio also affect performance. Similar observations apply to footwear on snow surfaces [126].
Finally, we note that measurements of rubber–ice friction on ice surfaces with a low concentration of snow crystals or with a thin snow film show drastically reduced friction coefficients [127]. Pilots have also reported that ice-covered runway surfaces are much more slippery when loose snow is present [106, 109].
It is not clear to us how these results can be understood, since ice sliding on ice generally gives relatively large friction coefficients, unless the temperature is close to the ice melting point or the sliding speed is so high that frictional heating becomes important [102, 128].

### v3

@@ 9 Rubber friction on ice and snow
Friction on ice is a fascinating and long-studied subject [101, 102, 103, 104, 105], with pioneering work by Faraday [106] over 150 years ago, who proposed that the surface of ice is covered by a liquid-like film of water. This phenomenon is now known as premelting. The presence of a premelted layer depends on both temperature and the hydrophobic or hydrophilic nature of the counter surface [107, 108]. For inert hydrophobic materials, such as polytetrafluoroethylene (Teflon), such a layer may form well below the melting point of ice, while for hydrophilic surfaces such as silica it may not occur at all, or only very close to the melting point.
The origin of this water film has long been debated. One explanation, proposed by Thomson [110], is pressure-induced melting, but this mechanism is unlikely since plastic deformation of ice usually occurs before the pressure is sufficient to induce melting. A more widely accepted explanation is frictional heating, suggested by Bowden and Hughes [111], who argued that sliding at sufficiently high velocity produces local melting of the ice surface. This view has been supported by experimental studies [112]. An important observation for friction on ice is that when the sliding speed increases the friction drops, which implies reduced frictional heating. This shifts the onset of ice melting to higher sliding speeds and can explain the velocity dependence of friction on ice [113].
More recent work has challenged the idea that interfacial water films alone are responsible for the low friction of ice. Molecular dynamics simulations by Atila et al. [109], inspired by the hypothesis of Moras et al. [114] on amorphization in incommensurate contacts, showed that ice surfaces form local “cold-welded” sites during sliding. These sites rapidly transform into amorphous layers under lateral displacement, without requiring significant heating or reaching melting conditions. The structure of the amorphous layer resembles supercooled water, providing a lubricating film that reduces shear stress. Its thickness scales with the square root of the sliding distance, indicating a displacement-driven process, while local pressure gradients further promote amorphization. In this framework, frictional heating plays a secondary role by lowering the effective viscosity of the amorphous layer, thereby reducing shear stress, while amorphization itself remains the dominant mechanism. The simulations further reveal that the mechanism operates similarly at hydrophilic and hydrophobic interfaces, although capillary and adhesion-enhanced dissipation increase friction for the former. This new perspective helps to explain observations that traditional theories could not account for, such as the persistence of interfacial water well below $0^{\circ}$C and the pronounced frictional contrast between hydrophilic and hydrophobic counterfaces. It also clarifies the dual temperature effect: at lower temperatures, amorphous layers form more rapidly but exhibit much higher viscosity, resulting in higher macroscopic friction.
@@
  • At temperatures close to the bulk melting point of ice, or at high sliding speeds, a water film or a low-viscosity amorphous layer is present at the interface. In this regime, the friction is dominated by the viscoelastic deformation of the rubber induced by the ice roughness [98, 107, 115, 109].
@@
  • At lower temperatures and lower sliding speeds, the interfacial water film is absent or the amorphous layer exhibits a much higher effective viscosity. In this case, the friction arises from both the viscoelastic response of the rubber and the shear in the rubber-ice contact regions, or between ice fragments attached to the rubber surface and the underlying ice, which corresponds to an ice-ice interaction [107, 109].
@@
  • At even lower temperatures (below $T_{\rm g}$), the viscoelastic contribution becomes negligible. As the rubber stiffens, it can cause wear of the ice asperities upon contact, leading to increased friction through a plowing-type process [115, 116].
@@
In Ref. [98], the Persson rubber friction theory was applied to rubber sliding on ice. The study incorporates both the adhesive contribution from the real area of contact and the viscoelastic contribution from the rubber. A temperature-dependent large wavenumber cutoff, $q_{1}(T)$, was introduced to account for the smoothing of short-wavelength roughness by plastic deformation of the ice. Since the ice penetration hardness depends on the temperature and indentation speed dependent, the cut-off depends on the ice surface temperature and on the rubber sliding speeds. The model also includes the effect of frictional heating in the rubber, which leads to a shift in the viscoelastic modulus to higher frequencies. The model offers a consistent explanation of the measured friction behavior for three very different tread compounds.
@@
Figure 63: (a) The measured rubber-ice friction coefficient as a function of the logarithm of the sliding speed for several indicated temperatures. At low temperatures, the friction is very similar to that of ice sliding on ice, suggesting that the friction originates from ice wear particles on the rubber surface sliding against the ice surface. (b) Comparison between experimentally measured friction coefficients (symbols) and theoretical predictions of the viscoelastic contribution (solid lines), based on the theory developed in Ref. [98]. At the higher temperature (black), the measured friction closely matches the viscoelastic prediction, indicating that a thin premelted water film forms at the rubber-ice interface and that friction is dominated by viscoelastic deformation. At the lower temperature (green), the measured friction is significantly higher than the theoretical prediction, suggesting additional contributions from interfacial shear. Adapted from Ref. [107].
@@
Figure 64: The friction coefficient as a function of the logarithm of the sliding speed for ice sliding on ice at different temperatures. Adapted from Ref. [117].
@@
Fig. 63(a) shows the rubber friction on ice as a function of the sliding speed for several different temperatures. In Fig. 63(b) the experimental measurements are compared with theoretical predictions based on the viscoelastic contact mechanics theory developed in [98]. At the higher temperature ($T=-5.4^{\circ}{\rm C}$), the experimental data nearly coincide with the predicted viscoelastic contribution, indicating that a thin interfacial film is present, which may consist of a premelted water layer or a low-viscosity amorphous layer. In this regime, the friction is dominated by viscoelastic deformation of the rubber induced by the ice roughness. In contrast, at the lower temperature ($T=-30^{\circ}{\rm C}$), the measured friction coefficient significantly exceeds the viscoelastic prediction. This suggests the absence of a liquid-like film or the presence of a highly viscous amorphous layer, so that friction arises not only from the viscoelastic response but also from interfacial shear in the rubber-ice contact regions. For $T<-20^{\circ}{\rm C}$ the friction of rubber on ice approaches that of ice on ice (see Fig. 64), indicating a substantial contribution from shear between ice fragments and the underlying ice surface.
@@
Figure 65: (a) A rubber block squeezed in contact with an ice surface covered by stone particles. (b) During sliding, ice plowing occurs, resulting in a friction force that is maximal when the projected area of the particles occupies the fraction $\psi\approx 2\sigma_{0}/\sigma_{\rm P}$ of the nominal surface area, where $\sigma_{0}$ is the nominal contact pressure squeezing the rubber block against the substrate, and $\sigma_{\rm P}$ is the ice penetration hardness (temperature dependent). The maximum plowing friction coefficient is $\mu\approx 1$. (c) The ice roughness with wavelengths larger than the sand particles contributes a viscoelastic component to the friction force.
@@
Suppose there are $N$ sand particles in the tire-road footprint. If the concentration of sand particles is high enough, the full normal load of the tire will act on the sand particles, giving the force $f_{\rm N}=F_{\rm N}/N$ that squeezes each sand particle against the ice surface. If the sand particles are spheres with radius $R$, and if $\sigma_{\rm P}$ is the ice penetration hardness, then the sand-ice contact area projected onto the horizontal $xy$-plane is $A=f_{\rm N}/\sigma_{\rm P}$.
The maximum friction occurs when the force $f_{\rm N}$ is large enough to squeeze half of the particle into the ice. If $f_{\rm N}$ exceeds the value required to embed half of the sand particle, the entire particle may be pressed into the ice, and the contact would become rubber-ice contact. For the stationary case, the contact area projected onto the horizontal $xy$-plane would be $\pi R^{2}$, and the maximum penetration resistance is $\pi R^{2}\sigma_{\rm P}$. However, for the sliding (plowing) case, we assume the contact area is half of this, $A=\pi R^{2}/2$, since the sand particle only contacts the ice on one side. In this case, the ice wear track will have a width of $2R$, and the vertical cross-sectional area of the plowing track will be $A_{\rm plow}=\pi R^{2}/2$.
@@
where $E_{*}=E/(1-\nu^{2})$. If this distance is larger than $R$, the rubber may contact the ice, and part of the normal force (the load) would then be carried by direct rubber-ice contact. As a result, the force $f_{\rm N}$ acting on the sand particles would be reduced. Hence, for (39) to be valid, we must have
@@
Figure 66: Measured rubber friction coefficient as a function of temperature for rubber sliding on clean ice, on ice with a thin snow film, and on ice covered by a thin snow film and sand. The friction was measured using the British Pendulum Tester (sliding speed on the order of $\sim 1\ {\rm m/s}$). Adapted from Ref. [121]
@@
Sanding with a high concentration of sand particles $\psi>0.2$ may result in friction smaller than that on a clean surface, unless the temperature is close to the ice melting temperature (see Fig. 66). This is consistent with some experimental observations, where the friction level of rubber on sanded ice was found to be lower than that on clean ice [122, 123]. Norwegian airport maintenance personnel have also pointed out that, in some cases, sanding has a negative effect on the friction level, and they commonly refer to it as a “ball-bearing” effect [121, 124]. (However, there is no general reason for the sand particles to roll on the ice surface.)
This behavior is expected when the sand particle concentration is high, because in this case (44) predicts $\mu<0.2$ if $\psi>9(\sigma_{0}/\sigma_{\rm P})$, and using $\sigma_{0}=0.2\ {\rm MPa}$ and $\sigma_{\rm P}=10\ {\rm MPa}$ gives $\psi>0.18$. Since sanding removes the adhesive contribution to friction and the viscoelastic contribution from roughness with wavelengths shorter than the sand particles, this may result in a friction coefficient smaller than that observed on clean ice for $T<-5^{\circ}{\rm C}$.
Finally, we note that ice wear tracks have been observed even when rubber slides on clean ice surfaces [125, 126]. In Ref. [125], it was suggested that the narrow wear tracks observed may have resulted from contamination particles (dust), which cannot be avoided when performing experiments in the normal atmosphere. However, in Ref. [126], it was suggested that hard filler particles on the rubber surface scratch the ice surface and result in the strongly modified (smoothed) ice surface observed after repeated sliding on the same area. In addition, the stress in rubber asperities that are strongly deformed (strain of order unity) is on the order of the modulus $E_{*}$, which can be similar to the yield stress of ice. Hence, wear tracks may arise if large enough rubber asperities occur, even without considering the fillers.
@@
The friction of rubber on snow shares many similarities with that on ice, as snow can be regarded as loosely packed ice particles. However, snowflakes can have a complex dendritic structure or consist of small, nearly spherical ice grains, and these two types of snow may exhibit different frictional properties against rubber or other solids. Although dedicated studies on the frictional behavior of rubber on snow surfaces remain limited (mostly for tires) [127, 128, 129, 130], there are many studies of the friction of polymer sliders (skis) on snow [131, 132, 133]. This field was pioneered by Bowden and Hughes [134], who used skis to investigate friction on snow and found characteristics similar to friction on ice, although they did not propose a specific model for snow friction. For stiff solid sliders such as skis, Colbeck [135, 136, 137, 138, 139] identified the key components of snow friction as plowing, direct solid-solid contact (i.e., adhesive friction), meltwater lubrication due to frictional heating, and capillary forces from liquid bridges in partially melted snow.
Some of these mechanisms for friction between snow and hard polymers may also apply to rubber materials, but additional considerations are required due to their viscoelastic nature (see Fig. 67). Any fluctuation in the height of the compacted snow layer that is not flattened in the tire footprint could generate a viscoelastic contribution to the friction during slip. However, if the frictional shear stress exceeds the shear yield strength of ice, the compacted snow may instead undergo internal shear deformations, which would also contribute to the friction force.
The shear strength of snow compressed by a normal stress $\sigma$ has been studied experimentally. Fig. 68 shows results for the shear stress immediately after compression (without aging), as relevant for tire applications. The shear stress increases nearly linearly with the applied pressure and is of similar magnitude to the normal stress. This implies that if the sheared area were similar in size to the nominal tire-road contact area, a friction coefficient of order unity could be expected. In reality, the area where the shear plane lies inside the compacted snow rather than at the rubber-snow interface may be much smaller than the nominal contact area. However, if the tire had a large concentration of cavities as in Fig. 69, and if the pressure between the snow and the rubber in the cavities were comparable to that in the nominal tire-road footprint elsewhere, then a large friction coefficient could be expected.
@@
Figure 69: If the normal pressure acting on the snow in the rectangular cavities of the rubber surface is similar to the nominal pressure acting in the tire-road footprint, then shearing the snow-snow interface as indicated by the dashed line could contribute to the friction coefficient. Depending on the slip velocity, this contribution could be of the order of the fraction of the nominal contact area covered by the cavities, e.g., 0.5 if the cavities cover $50\%$ of the total area.
Overall, the friction of rubber on snow depends not only on the material properties but also on the snow morphology and water content (which is temperature dependent):
  • Wet snow or slush (snow mixed with water, salt, mud, etc.) can act as a high-viscosity fluid that cannot be squeezed out of the tire-road footprint even at relatively low speeds. Hence, it may act as a high-viscosity lubricant, which could result in viscous hydroplaning and low friction.
  • On dry snow, friction arises in part from shearing snow contacts and in part from viscoelastic deformation of the rubber by inhomogeneities in the height of the compressed snow film. There is also a contribution from shearing the rubber-snow interface if slip occurs at this interface.
For rubber tires, tread design and sipes play a critical role. Compared to smooth surfaces, treads with sipes generate additional plowing at the edges of the bent rubber segments. In addition, compacted snow can enter the sipes, resulting in a snow-snow interface (of compacted snow) that is sheared during slip. These effects significantly enhance snow grip [127, 129, 130]. Tread depth and void ratio also affect performance. Comparable observations have been reported for footwear on snow surfaces [141]. Analogous to artificial tread structures, natural surface adaptations can also play a role in snow friction.
Recent experiments on bear paw pads have highlighted the importance of papillae morphology for frictional performance on snow [144]. Polar bears exhibit papillae that are significantly taller than those of brown and black bears, leading to a higher height-width ($H/D$) ratio. Friction tests with three-dimensional printed model surfaces demonstrated that shear stress on snow increases with increasing $H/D$, an effect comparable to the function of spikes in human footwear. As a result, polar bear paw pads achieve a 1.331.33-$1.50$ times higher shear stress compared with other bear species. Notably, polar bear paw pads also exhibit smoother roughness at short wavelengths. This is attributed to the presence of a thin water film at the ice-snow interface, which prevents roughness at these scales from penetrating the film and contributing to friction. Therefore, during evolution, polar bears did not develop pronounced short-wavelength roughness.
These observations emphasize the general principle that surface microgeometry, whether engineered in rubber treads or evolved in biological systems, can critically influence snow friction through its effect on interfacial shear stress.
Finally, we note that measurements of rubber-ice friction on ice surfaces with a low concentration of snow crystals or with a thin snow film show drastically reduced friction coefficients [142]. Pilots have also reported that ice-covered runway surfaces are much more slippery when loose snow is present [121, 124].
It is not clear how these results can be understood, since ice sliding on ice generally gives relatively large friction coefficients, unless the temperature is close to the ice melting point or the sliding speed is so high that frictional heating becomes important [117, 143].